How many moles of CF4 can be produced when 9.00 mol C react with 10.00 mol F2?

________ mol CF4

How many moles of each reactant are left over? Note that at least one must disappear.

______mol C remain

________ mol F2 remain
Answer #1

       C + 2 F2 ------------------- CF4

1 mole   2 mole

number of moles of C = 9.00 mol

number of moles of F2= 10.00 mol

according to equation

1 mole of C = 2 moles of F2

9.00 moles of C = ?

                                = 9.00 x 2/1 = 18.00 moles of F2

we need 18.00 moles of F2, but we have 10.00 moles of F2. So, F2 is completed first in the reaction.

Hence F2 is limiting reagent.

according to equation

2 moles of F2 = 1 mole of CF4

10.00 moles of F2 = ?

                                  = 10.00 x 1 / 2 = 5.00 moles

number of moles of CF4 formed = 5.00 moles.

according to equation

2 moles of F2 = 1 moles of C

10.00 moles of F2 = ?

                              = 10.00 x 1/2 = 5.00 moles of C

number of moles of C used in the reaction = 5.00 moles

remaining number of moles of C = 9.00 - 5.00 = 4.00 moles

number of moles of C remains = 4.00 moles

number of moles of F2 remains = 0.00 moles because it is completely involved in the reaction.